The SSSG group was very busy prior to sailing in order to accomodate specific requests from the GSO-Narragansett Broadscale group for a new exhaust hood/sink and for arrangement of the main lab benches. Hovey Clifford had little turn around time between OC297 and our leg to coordinate final design and construction of the fume hood, which, after some arbitration, completely satisfied our scientific and safety needs.
We thank and want to recognize a truly exceptional Captain and crew, and all those who made OC298 successful.
This cruise was sponsored by the National Science Foundation and National Oceanic and Atmospheric Administration. This report was edited by Maureen Taylor and Erich Horgan.
Phytoplankton, Chlorophyll, and Nutrients.....
Zooplankton and Ichthyoplankton Studies
Preliminary Results : Zooplankton
Preliminary Results : Ichthyoplankton
Preliminary Results : 10 m2 MOCNESS
Preliminary Results : Reeve net
Copepod Life History Studies
Appendix A. Personnel List
Appendix B. Event Log
Appendix C. Hydrographic Data
(1) determine the distribution and abundance of the ichthyoplankton and zooplankton community on the Bank and in the adjacent Gulf of Maine and slope waters. Emphasis is on target fish (eggs, larval and juvenile cod and haddock) and copepod species (all stages of Calanus finmarchicus and Pseudocalanus sp.) and their predators and prey.
(2) provide systematic collections of larval and juvenile cod and haddock for age and growth estimates and feeding habits.
(3) collect individuals of Calanus and the euphausid Meganyctiphanes norvegica, for population genetics studies.
(4) conduct a hydrographic survey of the Bank.
(5) map the Bank wide velocity field using an Acoustic Doppler Current Profiler (ADCP).
(6) deploy four shallow (10m) and one deep (40m) expendable drifter.
(7) to redeploy two long-term moorings which had become stationarily-challenged late last year.
locations which cover the entire bank (Figure 1A). The Broad Scale sampling protocol separates these 39 stations into two groups: full stations and partial stations. At the 19 full stations, a complete set of sampling operations is conducted. This involved a double-oblique bongo net tow, a Mark IV CTD cast, a 1-m2 MOCNESS (Multiple Opening Closing Net Environmental Sampling System, MOC-1) tow, a 10-m2 MOCNESS (MOC-10) tow, a plankton pump cast, and a vertical Reeve net tow. At the partial stations only the Bongo tow, CTD cast and MOC-1 tow are done. Seawater samples were collected during CTD operations for phytoplankton pigment and nutrient analyses. Additionally, samples for phytoplankton species identification, cell count, and spatial distribution were taken for later shore analysis. Occasionally, water was also drawn for salinity determination and H218O/H216O isotope concentration analysis. A Licor light meter was used opportunistically to measure Photosynthetically Active Radiation (PAR). At selected stations, the real-time CTD and a Niskin bottle cast were made for calibration purposes. Bongo tows are done in between the standard stations to increase the sampling density for cod and haddock larvae.
Five expendable drifters (4 shallow-10m, 1 deep-40m) were deployed at predetermined stations as per Dick Limeburner.
Current measurements were continuously collected by a hull mounted 300 kHz Acoustic Doppler Current Profiler (ADCP).
Two operations beyond the normal scope of the Broad Scale survey plan were also conducted during the cruise. First, two long-term moorings were redeployed at standard stations 9, on the Southeast Flank, and at standard station 20 on the Northeast Peak. Second, we agreed to recover another mooring which had broken free from its anchor.
After returning from sea, Peter Wiebe was apprised of the situation and returned a while later with news that the proper configuration files were present, but that the file for the conductivity sensor on the MOC1 was corrupt. This was the cause of our problems, not something inherent to the Windows 95 version of mocness.exe.
The Electroline termination used for the MOC10 would return from fishing the net with a half-turn on the bridle. The shackle holding the bridle to the termination was turned around a half turn after each haul. It is suspected that the new 0.68 wire on the winch
The Reeve net protocol needs to be modified. The total Reeve net weight on OC298 deployments was insufficient to maintain enough tension on the hydro winch wire during descent; the wire went slack as the ship rolled with the seas. We used a Bongo Ball attached to the termination to achieve proper wire tension. Additional weight is going to be added to the codend bucket bottom for the next Broad Scale cruise to assist in a proper wire tension during deployment.
There are four operations conducted from the hydro winch on Oceanus; we had to frequently break the "mecca" connections for each instrument: Seabird CTD, MK5 CTD, Pump, and 1-m MOCNESS. This resulted in a fair amount of stress on the mecca wires and we had to re-terminate adaptors on a several occasions. The CTD systems performed well during the entire survey. There were some minor problems with the lanyards on the MK5 rosette. It is now believed that the safety line above the instrument interferred with the bottles and caused occasional misfires.
Wednesday February 12. First station at approximately 01:50. Bongo and CTD operations went well. Problems with the 1 m2 MOCNESS at 05:30. E. Horgan discovered that the E/O plug was improperly attached to the step motor. The acquisition software would consistently display Error 5 and Error 11 messages upon start of program. Futile attempts to remedy the situation left us switching the CPU with Windows 95 on it to the backup one with Windows 3.1. This solution allowed us to fish the MOCNESS successfully. A second bongo haul was made in place of the 1 m2 MOCNESS at stn stations 1 and 2. MOCNESS (1m) arrangement not entirely satisfactory. The chinese finger (used for safety line attachment) comes very close to the block, so there is very little clearance off the deck when lifting the frame from its cradle. Also, it was agreed a tugger would be much safer and easier to use than a person when retrieving the frame from outboard. CTD systems are working well although some of the rosette bottles for water sample collection have mis-fired.
Thursday February 13. Have completed stn stations 4-7, Flow meter from 1 m2 MOCNESS is also being used on the 10 m2 MOCNESS because of a faulty reed switch. Otherwise, all other operations progressing smoothly.
Friday February 14. Arrive stn sta 009 at 07:15. Preparations for the 1st mooring deployment begin (Irish and O'Malley). Mooring successfully deployed at 11:25. A second MK5 CTD cast was made after the deployment. A light cast was made using the MK5 with a par sensor clamped above it (Townsend). The 10m2 MOCNESS haul was re-towed because of the safety line getting stuck in the flow meter. Winds and seas began to build by mid-afternoon. A second bongo haul was made in lieu of the 1 m2 MOCNESS at stn stations 10 and 11.
Saturday February 15. Stn stations 12 - 16 completed. Winds and seas again increasing, cod end #9 lost during 1 m2 MOCNESS. The MOC-10 was cancelled but the Reeve net was successfully deployed. The MOCNESS acquisition software displays a persistent "ERROR #5" which the SSSG technician, by process of elimination, determined to be from corrupted GPS input *. We have since disconnected that input and are again operational.
Sunday February 16. Stn sta #17, conditions still marginal for deploying MOCNESS systems. A second bongo haul was made in lieu of the 1 m2 MOCNESS. Plankton pump operations were aborted after the hose detached from the wire. All operations were successfully deployed (except 10m2 MOCNESS) at stn station #18. Because of the forecast for deteriorating weather during the next 36 hours, it was decided to skip stn sta #19 (and the intermediate bongos on either side of it) and head directly for stn station #20 and deploy the second Irish mooring. An attempt will be made to pick up stn #19 after exiting Canadian waters if time and weather permit. 1930 - Irish mooring deployed, station operations begin immediately following deployment.
* A post-cruise evaluation by Peter Wiebe revealed that the configuration file for one of the environmental sensors was corrupt and that, at least, contributed to the error messages we encountered.
Monday February 17. We began experiencing intermittent signal problems with the Seabird CTD on the bongo as well as the 1 m2 MOCNESS system. The problem was a bad wire on the termination. We swapped out 1 of the ground wires, repeated the bongo with no problems. However, at stn station #24 the 1 m2 MOCNESS system continued to lose signal during the haul. After three attempts, it was decided to perform a second bongo haul in lieu of the MOCNESS. The adapter that plugs the 1 m2 MOCNESS under water unit into the sea cable (e/o--> mecca) was found to be open. A quick repair was made, and the system would be tested at the next station.
Tuesday February 18. All operations went well at stn station #25: Bongo, Pump, CTD, 1 m2 MOCNESS, 10 m2 MOCNESS and second drifter deployment. Winds and seas once again building. During the Bongo haul at stn #39, a very large wave crashed over the side of the deck work area. The stands that hold the 1 m2 MOCNESS buckled and broke causing the frame to fall to the deck. The plankton pump barrel was also hit and badly damaged. The electronics were taken off the 1 m2 MOCNESS and tested in the lab. The system appears to be O.K. but the housings of the battery case and under water unit should be checked for any small fractures. The stainless steel bars that run along side of the frame were bent at the top and bottom threads and it is feared that there may be stress fractures. Weather decks were secured and operations suspended. We will not attempt to deploy the 1 m2 MOCNESS again since it is uncertain if the pressure housings of the electronics would leak and cause more damage, or if the frame itself has been weakened by the bent bars. The plankton crew will attempt to reshape the pump barrel.
Wednesday February 19. Station operations resume. Completed stn stations 26 - 28 and intermediate bongo stations.
Thursday February 20. Signal problems occur at stn #29 during the second bongo haul. We removed the 2:1 conductor adaptor (since the Seabird does not require two conductors) and the haul went well. A plankton pump (newly repaired) and Reeve net haul were then done to allow time to make up a new 2:1 adaptor (required on the MK5). There were no signal problems during the MK5 CTD cast although we were again plagued with mis-fired bottles on the rosette. Seas building 10 - 15ft, winds 40+ knots. We are monitoring the position of a loose GLOBEC mooring drifting in the Great South Channel and have been requested to assist in its recovery if time allows (R. Schlitz, NMFS). Our cruise track has been alterred to pick up Stn station #19 (skipped earlier in lieu of the 2nd Irish mooring). Station sequence will be 30,19,32,31,40,33,34,35,...38 A weighted tether was rigged by D. Mountain and J. Irish to aide in CTD deployments and prevent the MK5 from swinging during launch and recovery. This worked well on its first test. Winds and seas calming by 7:00 pm.
Friday February 21. At stn station #40, the 10 m2 MOCNESS was deployed first because of predicted increasing winds and seas (again!). Some signal problems with the Seabird during the first bongo haul. The screw that holds the ground wires to the termination was found to be loose. After tightening, the two bongo hauls and MK5 CTD cast went well. Still hoping to be able to assist in the recovery of the loose GLOBEC mooring in Great South Channel. Have requested radio contact with R. Schlitz to get the most recent position and complete trajectory since breaking free. Working at Stn station #34 @ 10:00 am. We should complete the cruise track by sunrise tomorrow and hope to spend the day searching for the wayward mooring. Radio call with R. Schlitz gave latest position 41.009 N 68.008 W @1100z. Pump at this station was repeated because of missing the "net change" cue from winch operator.
Saturday February 22. 10 m2 MOCNESS and Pump cancelled due to high winds and seas. Bongo and CTD operations continue. J. Irish repaired a broken mecca connector on the sea cable splice. Winds gusting over 40 knots. We lumbered our way to stn sta #38 at about 4 knots. Bongo, CTD, Reeve, and two drifters deployed. Last station completed. At 08:00 we receive a fax from R. Schlitz giving latest GPS fix on loose mooring. It is about 35 miles from our present location. Search pattern begins at 13:00 with at least six observers scanning the horizon for the yellow mooring. The buoy was spotted by Toni Chute at 15:00 (approximately 3 miles from the latest reported position). After a few passes, the mooring was skillfully brought on board by the Bosun and crew of the Oceanus at 16:50.
We are heading home with the forecast for a bumpy ride with winds over 45 knots.
Bongo hauls were made at each of the stations occupied. A Seabird Electronics Seacat model 19 profiling instrument (SBE19 Profiler) was used on each bongo tow to provide depth information during the tow. Pressure, temperature, and salinity observations are recorded twice per second by the Profiler.
The following is a list of the CTD data collected with each of the sampling systems used on the cruise:
Instrument # Casts MK5 40 MK5 calibration 38 SBE19/Bongo 103 SBE19 calibration 6
The MK5 was deployed with 11 bottles on the rosette and samples were collected for various investigators. At primary #1 and 2 stations, 400 mls were immediately siphoned out of two Niskins (bottom and mid-depth) for observations of micro- zooplankton swimming behaviour (S. Gallagher and H. Brown, WHOI). Samples were collected for oxygen isotope analysis at selected depths for R. Houghton (LDGO) and a sample was taken at the bottom for calibrating the instrument's conductivity data. Chlorophyll and nutrient samples were collected at various depths by D. Townsend and J. Xu. Surface samples for phytoplankton species composition were collected and preserved for J. O'Reilly (NMFS) at full stations only.
Parameter # samples taken Oxygen isotope 131 Micro-zooplankton 34 Chlorophyll 146 Species composition 12
The SBE19 Profiler and the MK5 data were post-processed at sea. The Profiler data were processed using the Seabird manufactured software: DATCNV, FILTER, ALIGNCTD, BINAVG, DERIVE, ASCIIOUT to produce 1 decibar averaged ascii files. The raw MK5 data files were processed using the manufacturer's software CTDPOST in order to identify bad data scans by "first differencing." The latter program flags any data where the difference between sequential scans of each variable exceed some preset limit. The "Smart Editor" within CTDPOST was then used to interpolate over the flagged values. The cleaned raw data were converted into pressure averaged, pressure centered 1 decibar files using algorithms provided by R. Millard of WHOI, which had been adapted for use with the MK5.
Both CTD systems work well. The Mecca termination connectors on the hydro wire failed on a few occasions due to the stress of changing between instruments on each station. The rosette suffered a number of misfires during the first part of the cruise, and the lanyards on some bottles hung up - resulting in some desired samples not being collected or being lost. A safety wire being used on the CTD package was removed because it appeared capable of interfering with the firing of the bottles, and the lanyards were rearranged to be positioned along the right side of the bottles. Misfires and lanyard hang ups largely disappeared and the second half of the cruise no samples were lost due to the earlier problems.
Due to weather, the MK5 system was not deployed on two stations (standard stations 30 and 39 - shown as station numbers 173 and 159, respectively, in Figure 1b), and the SBE19 data from the bongo tow at those stations were substituted as the primary hydrographic.
The volume average temperature and salinity of the upper 30 meters were calculated for the four sub-regions shown in Figure 8. These values are compared with characteristic values that have been calculated from the MARMAP data set for the same areas and calendar days. The volume of Georges Bank water (salinity < 34 psu) was also calculated and compared against the expected values.
The water temperature was near normal over most of the Bank, except along the southern flank, where the anomaly values were < - 1.0 c, particularly along the bottom (Figure 3b). The salinity was considerably lower than the MARMAP reference values, with anomalies over much of the bank of < -0.7, and over the southern flank, of < -1.0 (Figure 6). The lowest salinities observed were at the offbank stations along the southern flank and in Northeast Channel. These values suggest a westward flow of Scotian Shelf water across Northeast Channel and just south of the bank. The negative anomalies of bottom temperature and salinity suggest the influence of Scotian Shelf water over much of the southern flank of the Bank between 66 30' and 68 30' W. The Scotian Shelf water influence does not appear continuous from the Northeast Peak westward, with a break near 66 30' W. The Scotian Shelf water may have intruded across Northeast Channel in pulses, or may have moved west south of the Bank and intruded up onto the southern flank west of 66 30' W from the south. Analysis of satellite imagery may be able to resolve this uncertainty.
The large negative salinity anomalies over the Bank continue a trend toward lower salinity that has been observed since late 1994 - through the 1995 and 1996 GLOBEC sampling periods - in the Gulf of Maine and on Georges Bank. The intrusion of Scotian Shelf water along the southern edge and flank of the Bank is a separate (at least locally separate) phenomenon from the large negative salinity anomalies over the rest of the Bank.
The water columns on the Bank (< 80 m water depth) were well mixed, except on the eastern part of the bank, where the intrusion of Scotian Shelf water in the near surface layer (0-25m) resulted in modest stratification.
The fluorescence values (figure 7) are low relative to values observed later in the season in 1995 and 1996. Also, with the well-mixed conditions, there were no large subsurface peaks in fluorescence often observed in association with a well developed pyncocline later in the season. The highest fluorescence values during this cruise were in the shallow, central portion of the Bank. Figure 9 shows the comparison between the fluorescence data from the CTD (volts) and the extracted chlorophyll. The R2 is approximately 0.60.
The purpose of this project is to investigate the idea that the growth and production of zooplankton and fish on Georges Bank are limited by the amount of nutrients (especially nitrogen) that is brought onto the Bank from the nutrient-rich, deeper waters around the Bank's edges (cf. Townsend and Pettigrew, 1996). Thus, we are collecting water samples on four of the six broadscale cruises (February to May) to analyze for a suite of and phytoplankton biomass and chlorophyll a. The sampling period is chosen to bracket the winter-to-early summer transition. During this cruise, water samples were collected for the analyses of:
Water collections were made at various depths at all of the regular hydrographic stations (Stations 1 - 40) sampled during the February 1997 broad scale survey cruise aboard R/V Oceanus, using the 1.7 liter Niskin bottles mounted on the rosette sampler. Additional surface water samples were collected at positions between the regular stations (Stations 41 - 81) using a Kimmerer Bottle to sample a depth of 2m.
Light attenuation of photosynthetically active radiation (PAR) was measured at two stations, intended to be on the Bank at or about noon; these were the only such stations when sea state conditions allowed a light cast. A LiCor underwater spherical quantum sensor and deck-mounted cosine quantum sensor were used to compare the underwater light field as a function of depth and coincident surface irradiance. Data were collected at stations 9 and 18 (see Table 1). Samples for dissolved inorganic nutrients and chlorophyll were collected at all stations, 1-40, and all in-between stations (at 2 m). Water samples for DIN were filtered through 0.45 µm Millipore cellulose acetate membrane filters, and the samples frozen immediately in 20ml polyethylene scintillation vials by first placing the vials in a seawater-ice bath for about 10 minutes. Samples will be analyzed on shore follwing the cruise using a Technicon II 4-Channel AutoAnalyzer.
Water samples (50 mls) for dissolved organic nitrogen, and total dissolved phosphorus were collected at 2 depths (2 and 20m) at each of the main stations and frozen as described above. These samples will be analyzed ashore using a modification of the method of Valderrama (1981).
Samples for particulate organic carbon and nitrogen were collected by filtering 500 mls from 2 depths (2 and 20m) at each of the main stations onto pre-combusted, pre-ashed GF/F glass fiber filters, and filters frozen for analysis ashore. The filters will be fumed with HCl to remove inorganic carbon, and analyzed using a Control Equipment Model 240-XA CHN analyzer (Parsons et al., 1984).
Samples for particulate phosphorus will be collected as for PON (but 200 mls will be filtered) and frozen at sea. Laboratory analyses will involve digesting the sample in acidic persulfate and then analyzing for dissolved orthophosphate. Phytoplankton chlorophyll a and phaeopigments were measured on discrete water samples collected at all stations (see Table 2) and determined fluorometrically (Parsons et al., 1984). The extracted chlorophyll measurments involved collecting 100ml from all bottle samples taken at depths shallower than 60m, filtering through GF/F filters, and extracting the chlorophyll in 90% acetone in a freezer for at least 12 hours. The samples were analyzed at sea using a Turner Model 10 fluorometer. These data will be used in regressions against measurements of in situ fluorescence as part of the regular CTD casts.
Townsend, D.W. and N.R. Pettigrew. 1996. Nutrient limitation of secondary production on Georges Bank. J. Plankton Res. 19(2): In Press.
Valderrama, J.C. 1981. The simultaneous analysis of total nitrogen and total phosphorus in natural waters. Marine Chemistry 10: 109-122.
(1) Principle objectives of the ichthyoplankton group in the broadscale part of the U.S. GLOBEC Georges Bank Program were to study the composition of the larval fish community on Georges Bank, to define larval fish distribution across the Bank and within the water column, to determine those factors which influence their vertical distribution, and to determine bank-wide versus "Patch-Study" mortality and growth rates. Emphasis in this study is on cod and haddock larvae along with their predators and prey. This study also includes larval distribution and abundance, and age and growth determination. These objectives were implemented through use of bongo net and MOCNESS to make the animal collections.
(2) The primary objective of the zooplankton group was completion of a bank-wide survey of Georges Bank to determine the distribution, abundance, and stage composition of the target species Calanus finmarchicus and Pseudocalanus spp. A second objective was to identify, quantify, and describe the occurrence of abundant non-target species in order to provide a description of the environment occupied by the target species. These objectives were implemented by using the 1-m2 MOCNESS, a vertically discrete, multiple opening and closing net system for sampling the copepodite stages of the zooplankton, a submersible pump for sampling the naupliar stages, and a vertically towed Reeve net for sampling gelatinous predators.
In addition to these objectives, the zooplankton group was responsible for the following:
(a) To collect and sort Calanus finmarchicus at the fifth naupiar stage for RNA/DNA ratio analysis for Melissa Wagner at the University of Rhode Island.
(b) To take subsamples from the 1-m2 MOCNESS hauls for population genetic studies on Pseudocalanus spp. for Dr. A. Bucklin at the University of New Hampshire.
At stations where the 1-m2 MOCNESS system either was not towed or could not be used due to adverse weather conditions, a second bongo tow was made. This frame was fitted with both 335 µm mesh and 200 µm mesh nets. Digital flow meters were suspended in the mouth of each net to determine the volume of water filtered. Tows were made according to standard MARMAP procedures except maximum tow depth was 500 m. Wire payout and retrieval rates were ~40 m/min and 20 m/min respectively. Upon completion of the tow, the nets were each rinsed with seawater into a corresponding mesh sieve. The sample from the 200 µm mesh net was retained for zooplankton species composition, abundance and distribution, and preserved in 10% formalin. The other sample from the 335 µm mesh net was kept for molecular population genetic analysis of the copepod, C. finmarchicus, and preserved in 95% ethanol. After 24 h of initial preservation, the alcohol was changed. The used ethanol was retained for disposal or recycling ashore.
The 1-m2 MOCNESS sampler was loaded with ten nets. Nets 1-4 were fitted with 150 µm mesh for the collection of older and larger copepodite and adult stages of the zooplankton. Nets 0, and 5-9 were fitted with 335 µm mesh for zooplankton (nets 0 and 5) and ichthyoplankton (nets 6-9) collection. Tows were double oblique from the surface to within 5 m from the bottom. The maximum tow depth for nets 0, 1 and 5 was 500 m, and for net 6 was 200 m( if net 5 was sampled deeper than 200 m, it was returned up to 200 m and closed). Winch rates for nets 0-5 were 15 m/min and for nets 6-9, 10 m/min. The depth strata sampled were 0-15 m, 15-40 m, 40-100 m, and >100 ,m. The first (#0) and sixth (#5) nets were integrated hauls. For shallow stations, with only 2 or 3 of the depth strata, not all nets were fished. The contents of nets 0-4 were sieved through 150 µm mesh sieve, subsampled using a 2-L plankton sample splitter if the final volume was too large, then preserved in 10% formalin. Samples from nets 5-9 were sieved through 335 µm mesh sieve and preserved in 95% ethanol. After 24 h of initial preservation, the alcohol was changed. The used ethanol was retained for disposal or recycling ashore. At selected stations, 90-ml subsamples from the bottom and surface 150 µm mesh nets were removed and preserved separately in 10% formalin for Dr. C. Miller. At priority 1 and 2 stations, 90-ml subsamples from nets 2, 3, and 4 were removed and preserved in 95% ethanol. These samples were collected for Dr. A. Bucklin for population genetic studies to distinguish between Pseudocalanus species found on Georges Bank.
The 10-m2 MOCNESS was loaded with five 3.0 mm mesh nets. Tows were oblique from surface to ~10 m from bottom or a maximum depth of 500 m. The same depth strata were sampled as with the 1-m2 MOCNESS. The winch rate for retrieval varied between 5 and 15 m/min depending on the depth stratum. The slow winch rates were used in order to filter at least 4,000-5,000 m3 of water per depth stratum sampled. A stepped oblique tow profile during retrieval was used to achieve this, if needed. Catches were sieved through a 335 µm mesh, and preserved in 10% formalin.
The Pacer high-volume pump was used to collect nauplii and younger, smaller copepodite stages of zooplankton. The intake hose was deployed off the starboard side hydro boom by connecting the suction end, fitted with a 1.7-L Niskin bottle cut in half lengthwise, to the winch wire. The boom winch meter block was zeroed at the surface and the wire out reading was used to determine the depth of the cast. A 70 kg weight was used to depress the array. Three 30-m sections of 7 cm diameter hose were connected to the pump, allowing the intake hose to attain a maximum depth of approximately 85 m. At shallow stations, the intake hose nozzle was lowered to 3-5 meters off the bottom. Three integrated depth samples were collected with 35 µm mesh nets, sieved through a 30 µm mesh sieve and preserved in 10% formalin. Sampling depths were from the maximum depth to 36 m, 36-12 m, and from 12 m to surface. Before samples were collected, water was diverted from going into the net and was allowed to flush for 60 seconds. This assured that the zooplankton from the desired strata was obtained. Once at the surface, the intake section was held just below the surface for 60 s. This allowed the water to pass completely through the hose. Wire retrieval rate was approximately 4-5 m/min. This rate was used to obtain volumes of 500 L per 5 m depth interval sampled.
The Reeve net was used to collect larger gelatinous predators. The object is to try to sample larger gelatinous predators that are destroyed in the MOCNESS tows. The net was deployed off the starboard side hydro boom by connecting the net bridle to the winch wire along with a Bongo ball. The net was lowered at 15 m/min. Tows were made to the depth of the shallowest MOC-10 net, (usually 0-15 m). The winchman called out the wire out which was recorded by the science watch chief. The net was hauled up at about 5 m/min, and carefully retrieved over the side. The sample was then carefully poured through a large, stainless steel 303um mesh sieve, then resuspended in a liter jar and inspected for large gelatinous predators.
Protocol for collection of Calanus finmarchicus fifth naupliar stage for RNA/DNA ratio analysis include the addition of a fourth 30 µm mesh net during the regular pump cast. The pump intake was resubmerged after the completion of a normal haul to maximum depth or within five meters of the bottom. The RPM's of the pump motor was decreased by half in order to reduce the stress on the zooplankton. The rate of retrieval was approximately 3-4 m/min during collection, from bottom to surface. The animals caught in the net were poured directly into a bucket with seawater taken from the pump barrel. The sample was then concentrated into a 1 L plastic beaker and kept on ice. A small amount of sample was scooped into a petri dish and rinsed with MS-222 anesthetic solution and examined under dissecting microscopes. The intention was to fill five eppendorf tubes each with five N5 Calanus finmarchicus at standard stations 12, 20, 29, and 38. Upon examination of samples, however, we discovered a distinct lack of Calanus finmarchicus nauplii and were therefore unable to gather the number of specimens necessary for completion of this experiment.
Observations of nauplii collection at stations 12, 20, 29, and 38 reveals a low abundance of stage N5 Calanus finmarchicus. At stations 12, 20, and 29 abundances of all nauplii were low, with Pseudocalanus spp. being the most abundant species. Station 38 had the greatest abundance of C. finmarchicus nauplii, but still few N5's of this species were found. Observations of the 1-m2 MOCNESS samples from nets 0-4, and all pump samples will be staged and enumerated at the University of Rhode Island Graduate School of Oceanography GLOBEC Counting Laboratory.
No figures describing zooplankton species composition, distribution, and abundance are being provided in this report due to individual investigators variability in assessing the actual zooplankton picture when observing a small subsample of a complete haul. When the samples are fully processed and analyzed, the quantitative information collected in this cruise will become available and can then be used to provide a comparison of interannual variation of the biological community on Georges Bank.
1. Bongo nets, 0.61-m 102 tows 77 preserved,5% formalin 335 µm mesh 102 preserved,EtOH 200 µm mesh 25 preserved, 10%formalin 2. MOCNESS, 1-m2 15 tows 150 µm mesh 47 preserved, 10%formalin 335 µm mesh 15 preserved, 10% formalin 335 µm mesh 60 preserved, EtOH 3. MOCNESS, 10-m2 8 tows 34 preserved,10% formalin 3.0-mm mesh 4. Pump 14 profiles 41 preserved, 10% formalin 335 µm mesh 5. Reeve Net 11 hauls 11 preserved, 5% formalin 335 µm mesh
Cod (Gadus morhua)/Pollock (Pollachius virens): Cod/pollock larvae were collected consistently in low numbers (never more than five at a station; see cod/pollock figure 10), and showed a pattern of decreasing size as we sampled from the southwest toward the northeast portion of the Bank. As we rounded the Northeast Peak and began heading to the west, the size of the larvae observed in the samples gradually increased again. On the Northeast half of the Bank, the larvae observed were 5-6mm in length. On the Southwest half, their lengths ranged from 7 to 25 mm, averaging 12mm. The smallest larvae were found together with the greatest densities of cod/haddock/pollock eggs.
Atlantic herring (Clupea harengus): Catches of herring larvae were generally limited to the western portion of the Bank except for stations 70 and 31 on the north side and station 18 toward the east (see Atlantic herring figure 11). The highest concentrations of herring larvae were found in the southwest quadrant, the same as last month. Most larvae measured more than 30mm in length, and by next month will probably have grown out of range of our nets.
Sand lance (Ammodytes americanus): The ubiquitous sand lance larvae were col- lected at almost every station visited on this cruise (See figure 12). They were most abundant on the southwest area of the Bank, and concentrated in much the same area as the herring. In size, they ranged from 10 to 30mm; many catches contained several different sizes of larvae. The concentration of tiny larvae found in January, indicative of a localized spawning event, was not observed this month. The larvae appeared to be larger and more evenly distributed.
Cod/haddock/pollock eggs: Gadid eggs were concentrated on two areas of the bank; the Northeast Peak and the northwest quadrant (see figure 13). These areas also contained the highest densities of gadid eggs in January. The smallest cod/pollock larvae were found in conjunction with the greatest densities of eggs in the Northeast Peak area, although in low numbers. Next month, hopefully we will see an increase in numbers of larvae collected in these areas.
Standard Station 4, Haul 1
Atlantic herring (Clupea harengus) larvae, 1000 individuals averaging 30 mm Decapod shrimp (Crangon septemspinosa) Windowpane (Scopthalmus aquosus) larvae Standard Station 7, Haul 2 Euphausiids Ctenophores Large caridean shrimp Large chaetognaths Siphonophore bells Herring larvae Cod (Gadus morhua) larvae Standard Station 9, Haul 3 Ctenophores Naked pteropods Standard Station 13, Haul 4 Ctenophores Naked pteropods Cod larvae Hyperiid amphipods Siphonophores Standard Station 20, Haul 5 Ctenophores Euphausiids Hyperiid amphipods Decapod shrimp (Crangon) Standard Station 25, Haul 6 Ctenophores Euphausiids Hyperiid amphipods Gammarid amphipods Naked pteropods Siphonophores Standard Station 40, Haul 7 Hyperiid amphipods Euphausiids Ctenophores Standard Station 34, Haul 8 Large caridean shrimp Hyperiid amphipods Euphausiids Ctenophores
Renewed support for the OSU-based contribution to the GLOBEC Broadscale program is focused on a continuation of our examinations of life history patterns of Calanus finmarchicus on Georges Bank. Our projects fall into four major categories: (1) continued analysis of the frequency and environmental correlates of the apparent sex reversal of genetic male Calanus which occurs at maturation in a significant portion of the population, including correlation with fecundity data being gathered by Jeff Runge and proposed use of molecular methods to determine the genetic sex of individuals, (2) continued examination of jaw morphology as a diapause signature in fifth copepodites, and correlation with lipid storage and gonad development, and as an indicator of age-within-stage of all copepodite stages (3) analysis of fat storage by fifth copepodites, using images captured at sea and Charlie's new algorithm for calculating oil sac volumes in conjunction with gas chromatography, and (4) a substantial contribution to the GLOBEC modelling effort using "individual vector models", being developed by Charlie during his six month sabbatical in Nice, France.
Can Calanus males reverse their sex?
We have found definite seasonal trends in the proportions of quadritheks (possibly genetic males which have developed as females) in our Georges Bank samples from 1994, 1995 and 1996. We will continue to monitor this trend using formalin preserved subsamples from selected 150 micron MOC-1 and 200 micron bongo nets on this and subsequent Broadscale cruises. On this cruise, we collected subsamples (90/600ml) at standard stations 3, 4, 5, 7, 8, 9, 12, 13, 16, 18, 20, 23, 25, 39, 27, 29, 30, 40, 34, 36 and 38. This was a good month for sampling adults. Most of the Calanus found were females although there are still a number of males present.
We hope that we will be able to determine the underlying genetic sex of individual Calanus and correlate the quadrithek antennal morphology with genetic maleness. We are attacking this problem by analysis of DNA fragment lengths, which are expected to be different in X and Y chromosomes, and searching for highly multiple repeat sequences characteristic of sex chromosomes. For these analyses, we are cryopreserving adult male and female Calanus. On this cruise, adults were sorted from our subsamples and frozen live in liquid nitrogen from stations 5, 8, 9, 12, 18, 25, 29, and 38. Ethanol preserved subsamples from MOC-1 net 5 (90/400ml) taken at standard stations 3, 4, 5, 7, 8, 9, 12, 13, 16, 18, 20, 23, 25, 39, 27, 29, 30, 40, 34, 36 and 38 will also be used.
Jaw staging is also an indicator of an individual's age-within-stage. As the animal progresses through each stage, the jaw facies pass through recognizable as postmolt, late postmolt, intermolt and tooth formation phases. Preliminary analyses of jaw phases of individual second through fifth copepodites from 1995 Broadscale subsamples have yielded some interesting results with respect to the population dynamics of Calanus on Georges Bank. The formalin preserved subsamples listed above will be used for continuation of this effort.
Lipid analyses : total storage volume and component analyses We are studying the large store of oily wax which C. finmarchicus secretes into a tubular sac in the prosome of the fifth copepodite stage, prior to either maturation or rest. Actually, all copepodite stages have such sacs and accumulate some oil. The main question under study in 1997 is the areal and seasonal variation in quantities of oil in C5. Oil is quantified by an integration of oil sac projected area in video pictures and approximate conversion to oil volume, using image analysis and an algorithm recently worked out by Charlie for calculating an accurate volume estimate from the area. On OC298 there were very few C5's present in our subsamples, but sets of digital images were captured at standard stations 9, 25, 29 and 40. Fifth copepodites are recorded in groups of five, then cryopreserved for gas chromatographic analysis of the relative amounts of triacylglycerides and wax ester.
In early December 1996, problems were observed with the two Long-Term Moored Program's scientific moorings at the Northeast Peak and Southern Flank sites on Georges Bank. The Northeast Peak buoy had been showing a slowly decreasing battery voltage since deployment, and finally the solar power couldn't keep up with the power drain, and the system shut down due to low battery voltage on 13 December 1996. To that point the data telemetered and recorded internally was good. At that time the data from the Southern Flank mooring was checked more closely, and although the system was telemetering data regularly with full batteries, the data was not changing, indicating a problem with the main digitizer/sensor interface in the data system. Shiptime was obtained on the R/V OCEANUS with little advance notice when she came out of the ship yard and we were able to use some remaining GLOBEC shiptime. A quick recovery cruise was made from 18 to 20 December (R/V OCEANUS Cruise OC294) and both moorings were successfully recovered with no damage to the mooring components. Much thanks must go to the R/V OCEANUS and Captain Howland (on his last trip before retiring) for dedication to "getting the job done" in a professional and competent manner.
The buoys were returned to WHOI for servicing and repair. The Northeast Peak buoy (with the dead batteries) was disassembled, the data system powered up and tested. The current drain was normal and all indication were that there was no problems with the data system or sensors. The solar power regulator and batteries were then checked, and appeared to be working properly. Because the batteries were discharged down to about 6 volts (from standard 12 v), they were replaced with new ones, although it probably was not necessary as the old ones appeared to recharge and retain about their rated capacity. The solar panels were checked, and one was bad (no voltage output). This caused a loss of 1/4 of the charging capacity, but this alone should not have caused the problem. The other three solar panels appeared to be charging properly. Finally, the flashing guard light was checked, and it became apparent that here was the major problem. The bulbs were changed just before the cruise, and somehow bulbs with twice the current drain (1.1 amps at 12 v instead of 0.5 amps) were used, thus doubling the power. Then the flashing unit itself was drawing significantly more power than it should. Finally, the daylight shutoff switch was open, and not shutting off the light during the daylight hours. These things would marginally cause the observed loss of power, and were most probably the cause of the failure of the Northeast Peak mooring.
The Southern Flank mooring appeared to have a dead sensor interface unit. Upon opening and checking out the unit, it would not digitize reference signals or pass any diagnostic tests. So the EDIM (End Device Interface Module) was replaced with a new one fresh from calibration and the system appeared to be working properly again. There is the strong likelihood that the system was damaged by a static storm which was picked up on the signals leads of the meteorology sensors. This happened with the IMET suite of sensors on the SEWARD JOHNSON on SJ9504 during a snow squall in 1995, so is not that unusual. We had some problem with the wind sensor last winter, so there may be a basic problem that will have to be addressed in the future.
Good data was retrieved from both buoys up to the time of their failure, and this data was retrieved and is being processed. The new RDI Workhorse ADCPs and Sea Bird SeaCats were also checked and the data retrieved and processed. The WHOI bio-optical packages were also serviced, the data dumped, batteries replaced and the optical sensor windows cleaned. All sensors produced good data for the deployments. The systems were assembled with the telemetering temperature and conductivity sensors and run for a month on the WHOI dock successfully before deployment to test their operation.
The redeploymet was a problem as we did not have scheduled ship time, and no unscheduled time was available from local vessels. Ron Schlitz had a mooring deployment cruise in January, and we tried to get on that cruise, but with all their moorings gear there was no deck room for our equipment. Fortunately, we did not get on that cruise, as the weather only allowed their deployments and no further work of any kind after their moorings were deployed. We tried to get out the next weekend on last few days of their ship time which they did not need, but the wind was blowing 40+ kts and the seas at the Georges Bank NDBC buoy were up to 20 feet, so that opportunity was lost. The next opportunity was the two week GLOBEC Broadscale survey cruise on the R/V OCEANUS from 11 to 24 February 1997. Therefore, Jim Irish and Pat O'Malley packed up the buoys, sensors support gear, and jumped on the R/V OCEANUS for OC298.
The GLOBEC Broad Scale survey cruises work their way east along the Southern Flank, then west along the Northern flank making some 40 regular stations ( and many intermediate stations) with various sampling instruments and techniques used at each station. Two of their stations (Number 9 and 20) were at the Long-Term Moored Program's mooring sites. Friday 14 February the OCEANUS arrived at the Southern Flank Station. The buoy was swung down from the 01 level, and readied for deployment while standard broadscale survey activity continued to minimize the loss of time to the program. The same sensors were used in the same location as in the fall deployment (see cruise report R/V OCEANUS OC291). The Southern Flank mooring was ready for launch about 10:30 EST. The buoy was set in the water at 10:50 EST, the mooring was then paid out slowly by hand to the end of the sensor string. The ship had about 0.5 kts headway which was ideal and the deployment went very smoothly. Then the elastic tethers, and subsurface float were deployed so the mooring was strung out on the surface ready to tow into position at 11:12 EST. The ship slowly steamed into position between the two guard buoys and the anchor was dropped at 11:25 EST at 40 50.041' N x 67 19.209 W. Just a few tens of meters to the east and north of the previous position. The acoustic release was interrogated and disabled at 11:40 EST, at a range of 276 meters. The deployment went smoothly with the help of the ship's personnel, and the mooring appears to be deployed in the proper place and working. At 15:00 EST a radar check of the buoy indicated that it was still in the deployed position as three blips in a row were seen on the radar screen.
A supportive CTD profile was made by Dave Mountain near the mooring after deployment which showed a well mixed water column with a temperature of 4.9908 degrees C with a standard deviation of 2.2 milli degrees C. The salinity was also well mixed with a salinity of 32.3494 PSU with a standard deviation of 0.0005 PSU. The chlorophyll-a profiles and levels indicated that the spring bloom had not yet started. We may be able to observe it with the moorings this year.
On Sunday 16 February, the R/V OCEANUS skipped broadscale survey station 19 to arrive at the Northeast Peak site in daylight and good weather for the deployment of the mooring. The weather was sunny, cool and the wind was down to 10 kts. The seas are coming down also, but the weather forecast was for a full gale overnight, so a late afternoon deployment was selected to fit into the weather window.
The mooring was moved over from the port side, laid down on deck, and the sensors attached. The same sensors were used in the same location as in the fall deployment (see cruise report R/V OCEANUS OC291). This is the last deployment of this buoy system. It was first deployed in CODE as a guard buoy in 1980, then outfitted with a data system and successfully deployed in the Gulf of Maine in 1985-86 for 13 months. In this configuration, it was also deployed in Massachusetts Bay in 1989-90, and at the GLOBEC Crest site in 1994-95. Each of these deployments was for more than one year. This mooring had sensors at 10 meter increments to 50 meters, a downward looking ADCP at 3 m and a bio-optical package at 10 meters. The buoy was set up in 1.5 hours including moving the buoy and laying it down on deck. The mooring was deployed immediately after supper. The buoy was picked up at about 19:10 EST and the anchor splashed down at 19:26 EST. The position of deployment was 41 43.962 N x 66 32.152' W, about 0.04 nm from the previous position. The acoustic release was disabled and mooring operations shut down for the rest of the cruise.
A CTD profile was made by Dave Mountain near the mooring as an in situ calibration. This profile showed slightly increasing temperature profile (mean of 4.7999 degrees C with a standard deviation of 0.0159 degrees C and a warming with depth of about 0.05 degrees C). The salinity also slightly increased with depth (mean of 32.305 PSU, standard deviation of 0.0032 PSU and increase of 0.011 PSU). Much of the Bank has been well mixed to depth, but with horizontal gradients between stations. Therefore, there is something different going on in the Northeast Peak region. The ships drift before deployment was about 1.5 kts, which is slightly slower than observed in the previous deployment cruise.
The GLOBEC Long Term Moored effort would like to thank the Broad Scale Survey Crew (especially Erich Horgan, Maureen Taylor, David Mountain and Jaimie Pierson) and Captain Larry Bearse, Mate Courtney Barber, and Bos'n Jeff Stolp and the deck crew (Horace Medeiros, Chris Grimes, and Dave Philbrick) of the R/V OCEANUS for making this successful deployment of the moorings possible.
The crew then broke out the deck in preparation for recovery under the direction of Bos'n Jeff Stolp and the recovery started about 1350 EST. Because the weather was rough, extra care was made to set up all the lines, air tuggers, capstan, etc. before recovery started to speed up the operation as much as possible for safety. There was some difficulty picking up the buoy since the wind and currents were not favorable, but finally it was hooked, quickly brought aboard and secured. The VMCM was recovered over the rail next and taken out of the mooring line. The mooring tackle was then transferred to a line over the stern and pulled in on the capstan. The recovery went smoothly and an excellent job was done by the ships crew assisted by some of the scientific party.
The mooring hardware was all there, and appeared to be in good shape. One shackle in the 1/2" shot of chain below the current meter had the cotter pin almost out (it had never been bent over). The entire mooring was recovered down to the sling link at the anchor. Therefore, the mooring failed at the shackle to the anchor!. The current meter was relatively clean, and the fans were spinning freely. The guard light had some condensation on the inside of the lens, and as it got dark, it did not flash. The ARGOS transmitter was monitored by a Telonics Uplink receiver and we did not receive signals (as was the spare Limeburner drifter on the after deck). However, since the fixes were obtained by ARGOS, it is assumed that they was just a conflict of transmit times, but the unit should be checked. Everything was lashed down for transit to WHOI and at 1715 EST the OCEANUS headed for home.
NAME RATING 1. Lawrence T. Bearse Master 2. Courtenay Barber III Chief Mate 3. Anthony D. Mello 2nd Mate 4. Jeffrey M. Stolp Boatswain 5. David Philbrick OS 6. Horace M. Medeiros AB 7. Christopher Griner AB 8. Richard W. Smith Jr. Engineer 9. Alberto Collasius Jr. Jr. Engineer 10. Richard Morris Chief Engineer 11. Hugh B. Dakers Steward 12. Jovinol J. Fernandes, Jr. M/A
SCIENCE PARTY13. Erich F. Horgan (Chief Scientist) WHOI, Woods Hole,MA 14. Maureen Taylor (Co-Chief Scientist) NMFS, Woods Hole,MA 15. John Sibunka NMFS, Sandy Hook, NJ 16. Alyse Weiner NMFS, Sandy Hook, NJ 17. Antonie Chute NMFS, Narragansett, RI 18. David Mountain NMFS, Woods Hole, MA 19. James Gibson URI, Narragansett, RI 20. Maria Pilar Heredia URI, Narragansett, RI 21. Alyce Jacquet URI, Narragansett, RI 22. James Pierson URI, Narragansett, RI 23. Peter Clarke Mass. Maritime Academy 24. David Townsend U of Maine, Orono 25. Jiandong Xu U of Maine, Orono 26. Jennifer Crain OSU, Corvallis, OR 27. Pamela Arnofsky WHOI, Woods Hole, MA 28. James Irish WHOI, Woods Hole, MA 29. Patrick O'Malley WHOI, Woods Hole, MA 30. Heather Brown Volunteer 31. Laura G. Stein SSSG Technician